### v2

# A Decade-long Landscape of Advanced Persistent Threats: Longitudinal Analysis and Global Trends

Abstract.
An advanced persistent threat (APT) refers to a covert and long-term cyberattack, typically conducted by state-sponsored actors, targeting critical sectors and often remaining undetected for long periods. In response, collective intelligence from around the globe collaborates to identify and trace surreptitious activities, generating substantial documentation on APT campaigns publicly available on the web. While a multitude of prior works predominantly focus on specific aspects of APT cases, such as detection, evaluation, cyber threat intelligence, and dataset creation, limited attention has been devoted to revisiting and investigating these scattered dossiers in a longitudinal manner.
The objective of our study lies in filling the gap by offering a macro perspective, connecting key insights and global trends in the past APT attacks. We systematically analyze six reliable sources— three focused on technical reports and another three on threat actors— examining 1,509 APT dossiers (i.e., totaling 24,215 pages) spanning from 2014 to 2023 (a decade), and identifying 603 unique APT groups in the world. To efficiently unearth relevant information, we employ a hybrid methodology that combines rule-based information retrieval with large-language-model-based search techniques. Our longitudinal analysis reveals shifts in threat actor activities, global attack vectors, changes in targeted sectors, and the relationships between cyberattacks and significant events, such as elections or wars, which provides insights into historical patterns in APT evolution. Over the past decade, 154 countries have been affected, primarily using malicious documents and spear phishing as the dominant initial infiltration vectors, and a noticeable decline in zero-day exploitation since 2016. Furthermore, we present our findings through interactive visualization tools, such as an APT map or a flow diagram, to facilitate intuitive understanding of the global patterns and trends in APT activities.
Advanced Persistent Threats; Longitudinal Analysis; Global Trends
journalyear: 2025copyright: ccconference: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security; October 13–17, 2025; Taipei, Taiwan.booktitle: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security (CCS ’25), October 13–17, 2025, Taipei, Taiwandoi: 10.1145/3719027.3765085isbn: 979-8-4007-1525-9/2025/10ccs: General and reference Measurementccs: Security and privacy Social aspects of security and privacy
1. Introduction
Advanced persistent threats (APTs) are covert and sophisticated cyberattacks, typically orchestrated by state actors. APT campaigns attempt to gain unauthorized access to remote machines and stay undetected for extended periods, enabling targeted campaigns against governments and financial institutions (Amit Sharma and Saraswat, 2023). Their primary objectives are to steal sensitive data, to disrupt critical operations, and to undermine national security or economic stability (Adel Alshamrani and Huang, 2019).
Given the severity of risks and threats posed by APTs, both industry and academia put in constant effort to monitor and understand APT-involving incidents. The security industry (e.g., companies, experts, practitioners) investigates each APT incident, producing invaluable technical reports and articles. These scattered dossiers from varying sources provide detailed information on attack strategies, threat actors, and exploitation techniques on the web. Moreover, security practitioners track technical reports (CyberMonitor, 2024; Bandla and Castro, 2025; FKIE, 2024) and threat actors (MISP, 2025; StrangerealIntel, 2023; Cristaldi, 2025), organizing them in (public) repositories. The security industry also compiles Cyber Threat Intelligence (CTI) (Wikipedia, 2025d) databases to detect, analyze, and mitigate threats, including 1 Indicators of Compromise (IoCs) (Fortinet, 2025) for collecting forensic evidence, 2 Common Vulnerabilities and Exposures (CVEs) (Corporation, 2024a) for maintaining publicly disclosed security flaws, 3 MITRE ATT&CK Techniques (Corporation, 2024b) for identifying adversarial tactics, and 4 YARA rules (VirusTotal, 2022) for detecting specific malware patterns. For the intuitive understanding of APTs, visual representations with a map (Lab, 2025b; Kuhnert, 2025; Cristaldi, 2025) have been implemented on the web. Orthogonally to the aforementioned work by the cyber-security industry, academics also invest resources in better understanding past APTs. The majority of recent academic studies have focused on studying and analyzing specific attributes related to APT incidents, such as detection and evaluation (Aaron Zimba and Chishimba, 2020; Enes Altinisik, 2023; Wang et al., 2025; Mirco Marchetti and Guido, 2016; Milajerdi et al., 2019; Jia et al., 2024; Wang et al., 2024; Vaibhav Malik and Nalluri, 2024; Wajih Ul Hassan and Marino, 2020; Shen et al., 2024), CTI (Siracusano et al., 2023; Martin Ussath and Meinel, 2016; Nanda Rani and Shukla, 2023; Bahrami et al., 2019; Kumarasingh et al., 2024; Lv et al., 2024), and APT dataset (Siracusano et al., 2023; Lv et al., 2024; Bahrami et al., 2019; Kumarasingh et al., 2024; Branka Stojanović and Kleb, 2020).
In this paper, we observe that few studies have systematically investigated the landscape of APTs over an extended period of time. To help draw a complete picture of APT activity, we study the landscape of APT incidents over a period of 10 years. We uncover longitudinal changes through an in-depth analysis of fragmented, publicly available APT documentation. Our analysis can assist in identifying broader trends and patterns in surreptitious APT activities, offering valuable insights into the evolution of APT targets, malware samples, and sophisticated attack techniques.
More specifically, this study attempts to identify global trends from a macro viewpoint, including vulnerability exploitations, threat actor behaviors, and target changes through a large-scale investigation of $1,509$ unique APT dossiers and $603$ APT groups over the past decade. Due to the large volume of publicly available APT reports (i.e., $24,215$ pages of technical reports), we adopt a hybrid information-retrieval approach by leveraging the inference capabilities of Large Language Models (LLMs) combined with a rule-based extraction tool. To boost the LLM’s accuracy, we carefully design the questions and prompts, evaluating multiple models to identify the most effective retrieval for our goals.
With a comprehensive collection of APT dossiers (CyberMonitor, 2024; Bandla and Castro, 2025; FKIE, 2024), our study aims to analyze 1 the evolution of APTs over the past decade in terms of victim countries, threat actors, target sectors, initial attack vectors, and zero-day vulnerabilities; 2 CTI records in APT cases (e.g., CVEs, MITRE ATT&CK, YARA rules); 3 common traits of APT campaigns that demonstrate concealment (e.g., attack duration) and aggressiveness; and 4 external factors affecting APT campaigns, such as political events, international conflicts, global pandemics, or economic instabilities.
Our analysis yielded the following findings: 1 Over the past decade, APT campaigns have impacted 154 countries (80% of all nations), with the United States, India, and South Korea among the most frequently targeted. While $446$ unique threat actors have been identified in our dataset, a small set of actors is responsible for a significant share of attacks, including Lazarus (Wikipedia, 2025g), APT28 (Wikipedia, 2025f), and APT29 (Wikipedia, 2025c). In terms of targets, the government and corporate sector attract the majority of APT attention, with malicious documents and spear phishing serving as the dominant initial infiltration vectors. 2 Vulnerability-wise, while the exploited CVEs are highly severe (average score of $8.5$), our findings indicate that many of the attacks do not need to rely on zero-day vulnerabilities to be successful, which peaked between 2014 and 2016 but has declined thereafter. 3 In terms of the lifetime of the recorded incidents, APT campaign duration varies widely, from a single day to nearly five years (137 days on average). 4 Finally, an important finding is that APT activity frequently coincided with political events, international conflicts, global crises like COVID-19, indicating that attackers had already performed target reconnaissance and were waiting for an opportune time to act upon their findings.
To facilitate the exploration of APT campaign data by reviewers and eventually the general public, we designed an interactive map  [^1] https://lngt-apt-study-map.vercel.app/ that incorporates decade-long historical data, including threat actor(s), CVEs, attack vector(s), malware, target sector(s), and estimated duration, with support for dynamic updates using LLMs to retrieve content from technical reports. Additionally, we provide a flow diagram  [^2] https://public.tableau.com/views/TopMentionedCountries/Top30Countries illustrating the relationships between selected threat actors and target countries.
In summary, this paper makes the following original contributions:
  • We conduct a longitudinal measurement study of APT campaigns over the last decade (2014 – 2023), organizing $1,509$ unique technical reports, $603$ threat actors, and $177$ news articles.
  • To unveil longitudinal APT campaigns, we retrieve and refine responses to 10 identified questions using (context-aware) LLMs.
  • We carefully define four research questions centered on the evolution, CTI records, common traits, and external dynamics of APT campaigns over a decade. Our findings reveal global trends and key insights, including that the campaigns have affected 80% of countries worldwide; a small number of actors are responsible for a disproportionate share of attacks; and the exploitation of both zero-day and one-day vulnerabilities is prevalent.
  • We publicly release our curated dataset  [^3] https://zenodo.org/records/16869733 and an interactive APT-campaign map to foster future research in the field of APT studies.
2. Background
This section provides some background on advanced persistent threats, CTI, and the focus of prior work on APT campaigns.
Table 1. Summary of prior work on APTs. We classify them into five categories where the majority of those studies focus on specific domains, covering limited periods. Our work aims to offer insights into the evolution of APTs over the past decade from macro perspective (Section 2).
Category Topic Focus
Survey APT Survey (Adel Alshamrani and Huang, 2019; Amit Sharma and Saraswat, 2023)
Detection and evaluation Detection techniques (Aaron Zimba and Chishimba, 2020; Enes Altinisik, 2023; Wang et al., 2025; Mirco Marchetti and Guido, 2016; Milajerdi et al., 2019; Jia et al., 2024; Wang et al., 2024; Vaibhav Malik and Nalluri, 2024) Evaluating APT detection systems (Vaibhav Malik and Nalluri, 2024) APT reconstruction (Wajih Ul Hassan and Marino, 2020; Shen et al., 2024)
Cyber Threat Intelligence Information retrieval (Siracusano et al., 2023; Martin Ussath and Meinel, 2016; Nanda Rani and Shukla, 2023; Bahrami et al., 2019) Information recognition (Kumarasingh et al., 2024; Lv et al., 2024)
APT Dataset Dataset creation (Siracusano et al., 2023; Lv et al., 2024; Bahrami et al., 2019; Kumarasingh et al., 2024) Dataset evaluation (Branka Stojanović and Kleb, 2020)
Technical articles APT trends (GREAT, 2024; Traynor, 2024) Special reports (CISA, 2020; Jenkins and Black, 2024)
Advanced Persistent Threats. APTs refer to a critical and insidious category of cyberattacks characterized by their sustained, targeted, and highly sophisticated nature. Typical attacks include individual actors attacking any and all systems for financial gain, activism, or to demonstrate technological proficiency. However, APTs are often executed by well-organized, resource-rich groups with strategic objectives (Adel Alshamrani and Huang, 2019). These entities persist in their efforts to infiltrate and maintain a presence within a network over extended periods, often evading detection. Evidenced by exploiting zero-day vulnerabilities, deploying custom-developed malware, and adopting advanced social-engineering techniques, APTs’ resource-intensive operations indicate clear underlying motives - political (Jenkins and Black, 2024; Yerushalmy, 2024; Dearden, 2017), economic (Team, 2024; Legezo, 2016; Paganini, 2020), and military (Rogan, 2024; Zawya, 2024; of the Treasury, 2024). APT campaigns predominately aim at espionage, disruption, or sabotage, posing serious threats to national security and critical infrastructure.
Cyber Threat Intelligence. Cyber Threat Intelligence (Wikipedia, 2025d) is an essential pillar of modern cybersecurity for threat detection, prevention, and response. CTI enables organizations to collect and analyze information on cyber threats, anticipate emerging threats, and mitigate potential risks beforehand, empowering them to take proactive measures against cyberattacks. Indicators of Compromises (Fortinet, 2025) are critical elements of CTI, representing specific pieces of evidence that signal malicious activities within a system or network. Integrating such artifacts with CTI serves as actionable data for detecting and preventing future threats. IoCs are often aligned with standardized CTI frameworks to maximize their effectiveness. For instance, the MITRE ATT&CK (Corporation, 2024b) framework categorizes adversarial behaviors into distinct techniques, offering a structured approach to understanding how IoCs align within the attack lifecycle. In a similar vein, CVEs (Corporation, 2024a) maintain up-to-date (publicly known) vulnerabilities exploited by attackers, while YARA rules (VirusTotal, 2022) facilitate the detection of malicious artifacts through signature-based matching. As such, various tools have been developed to automate the extraction of IoCs from technical reports. One notable example is IoCParser (Tines, 2025), which specializes in retrieving diverse IoCs from URLs and texts. We leverage IoCParser to extract three primary IoCs – CVEs, MITRE ATT&CK technique IDs, and YARA rules.
Previous APT Studies. section 2 summarizes prior research on APTs. While extensive surveys on APTs have been conducted (Adel Alshamrani and Huang, 2019; Amit Sharma and Saraswat, 2023), the majority of these studies are dedicated to specific domains or cover limited periods. Beyond those surveys, we classify existing work into four distinct areas. First, a significant portion of the literature concentrates on detection and evaluation by exploring detection techniques (Aaron Zimba and Chishimba, 2020; Enes Altinisik, 2023; Wang et al., 2025; Mirco Marchetti and Guido, 2016; Milajerdi et al., 2019; Jia et al., 2024; Wang et al., 2024; Vaibhav Malik and Nalluri, 2024), evaluating existing detection systems (Vaibhav Malik and Nalluri, 2024), or reconstructing APT scenarios (Wajih Ul Hassan and Marino, 2020; Shen et al., 2024). Second, several studies focus on information retrieval (Siracusano et al., 2023; Martin Ussath and Meinel, 2016; Nanda Rani and Shukla, 2023; Bahrami et al., 2019) and information recognition (Kumarasingh et al., 2024; Lv et al., 2024) contributing to a deeper understanding of Cyber Threat Intelligence. Third, another direction explores APT datasets in terms of their creation (Lv et al., 2024; Bahrami et al., 2019; Kumarasingh et al., 2024) and evaluation (Branka Stojanović and Kleb, 2020). Lastly, the security industry has analyzed APT trends (GREAT, 2024; Traynor, 2024) and issued special reports on particular incidents (CISA, 2020; Jenkins and Black, 2024). However, our work differs from prior work by offering a longitudinal analysis based on a multitude of scattered APT dossiers, providing valuable insights into the evolution of APTs over time.
Figure 1. Overview of our methodology for longitudinal APT analysis. We collect technical reports, threat actors, and news articles across the web (Section 3.1). Then, we probe valuable information from technical reports based on rules and LLMs (Section 3.2). Note that we manually inspect attack duration (*) for precise analysis. Next, we refine raw information via normalization, categorization, de-duplication, and filtering (Section 3.3). Lastly, we conduct in-depth analyses to answer our research questions (Section 4).
Threat Maps. Geo-location maps are often used to visualize APT incidents. Kaspersky (Lab, 2025b), for example, offers a cyber attack map that depicts real-time worldwide cyber assaults. However, this tool needs to include historical data for APT activities. Both APT threat actor map (Kuhnert, 2025) and APTMAP (Cristaldi, 2025) aim to map APT-related data. However, both focus on APT groups’ physical locations, which are missing a longitudinal study. Although these works introduce an APT map, they focus on the distribution of victim nations by year alone. In this work, we devise a new APT map that represents combined information associated with a certain threat actor or victim country (Figure 10 in Appendix).
3. Methodology
This section sketches our methodology to reveal the landscape of APT campaigns over the past 10 years, including details on source collection, information retrieval, and raw data refinement.
Overview. Figure 1 illustrates the overview of our methodology to analyze decade-long APT cases spanning from $2014$ to $2023$. We explore open APT technical reports that provide details of an individual attack, threat actors, and varying news articles related to APT attacks (Section 3.1). Among those sources, we choose the dossiers that are 1 publicly accessible, 2 offering up-to-date information, and 3 written by a trustworthy entity. Additionally, we gather past articles from an authoritative outlet specializing in security-focused news. Considering the volume of technical reports (i.e., $24,215$ pages), we retrieve useful information (e.g., victim countries, target sectors, attack vectors) by leveraging an LLM that helps contextual inference. Note that we also utilize a rule-based (i.e., regular expressions) tool that performs better than the LLM probe when looking for specific information, such as, MITRE IDs, CVEs, and YARA rules. (Section 3.2). Subsequently, we refine the raw data using normalization, categorization, deduplication, and filtering (Section 3.3).
3.1. APT Source Collection
Table 2. Statistics on our collection of technical reports (TRs) and news articles on APTs. Out of 2,563 TRs, we analyze $1,509$ unique TRs (after removing $1,003$ duplicates and $51$ APT trend dossiers), along with $177$ news articles. The numbers in parentheses indicate the sum of all TRs before refinement. We discuss the credibility of TR sources in Section 3.1.
Year TR#1 (CyberMonitor, 2024) TR#2 (Bandla and Castro, 2025) TR#3 (FKIE, 2024) All TRs News Articles
2014 128 104 16 124 (248) 2
2015 150 87 28 135 (265) 8
2016 171 104 13 168 (288) 14
2017 124 87 26 142 (237) 15
2018 169 24 46 160 (239) 18
2019 222 27 57 201 (306) 15
2020 226 10 105 207 (341) 21
2021 156 13 94 162 (263) 19
2022 51 77 117 136 (245) 30
2023 21 42 68 74 (131) 35
Total 1,418 575 570 1,509 (2,563) 177
As a primary source, we use a series of open datasets that collect technical reports and threat actors. We acknowledge the collective intelligence of security experts and practitioners who collaborate to uncover covert threat activities around the globe. Of all, we carefully select reliable sources, which focus on three collections of technical reports (CyberMonitor, 2024; Bandla and Castro, 2025; FKIE, 2024) (subsection 3.1) and another three collections of threat actors (MISP, 2025; StrangerealIntel, 2023; Cristaldi, 2025) (subsection 3.1). While the former provide in-depth analyses of each APT incident, the latter offer individual APT group information. We include both sources because technical reports often lack detailed information about APT groups.
Table 3. Statistics on our collection of threat actors (TAs). Out of 1,684 TAs, we organize 603 unique TAs after removing 800 duplicates and 281 entries containing no information beyond their names. Note that the information of APT groups has been maintained by reliable sources (Section 3.1).
Collection Source of Threat Actors’ Information # of TAs
TA#1 (MISP, 2025) MISP Project (Cynthia Wagner and Iklody, 2016) 562
TA#2 (StrangerealIntel, 2023) Palo Alto (Networks, 2025), IBM X-Force(Security, 2025a, b), Malpedia (Daniel Plohmann and Padilla, 2017), Kaspersky (Lab, 2025c), Crowdstrike (CrowdStrike, 2025), Mandiant (Mandiant, 2025), Secureworks (Secureworks, 2025), Dragos (Dragos, 2025), Venafi (Venafi, 2025), CERT-UA (CERT-UA, 2025), Microsoft (Microsoft, 2025a) 692
TA#3 (Cristaldi, 2025) MITRE ATT&CK (Corporation, 2024b), ETDA ((2019), ETDA), VX-underground (Vx-Underground, 2025) 430
Total 603 (1,684)
Collection of Technical Reports. We thoroughly examined the available web sources to obtain the most reputable technical reports (TRs), which we refer to as dossiers. We adopt three TRs as reliable sources in this work. First, TR#1 (CyberMonitor, 2024) is a GitHub repository containing an extensive collection of $1,418$ TRs. This repository is continually updated as new dossiers on APT cases become available. Similarly, TR#2 (Bandla and Castro, 2025) is another GitHub repository that hosts $575$ TRs documenting various APT campaigns. Lastly, TR#3 (FKIE, 2024) is a curated list of $570$ TRs included as part of the comprehensive dataset maintained by Malpedia (Daniel Plohmann and Padilla, 2017). As shown in subsection 3.1, we aggregate these three sources as the collection of $1,509$ technical reports between 2014 and 2023 for our longitudinal analysis of APT cases. Although a small number of old TRs are available, we define our collection from the year of 2014 that shows a significant rise in the volume of TRs. Similarly, we exclude 2024 due to the small number of available TRs. Finally, we convert all web page reports into the PDF format for consistency.
Figure 2. Top 15 sources from the collection of technical reports. Most reports come from reputable sources such as Kaspersky (Lab, 2025a) and Trend Micro (Micro, 2014). We confirmed that 1,412 (93.6%) TRs are highly credible (Section 3.1).
Collection of Threat Actors (APT Groups). We separately collect threat actors (TAs) since the collection of TRs does not contain that information. First, TA#1 (MISP, 2025) includes information about the threat actors obtained from the MISP project (Cynthia Wagner and Iklody, 2016), one of the well-known Open Source Threat Intelligence Platforms. Second, TA#2 (StrangerealIntel, 2023) is maintained by diverse security companies and non-profit organizations which includes a wide range of APT groups. Third, TA#3 (Cristaldi, 2025) updates APT group information independently from the above two sources. Similar to the collection of TRs, we combine all the above three sources, obtaining the collection of $603$ unique threat actors. subsection 3.1 provides the statistics of our collection, which encompasses the information on TAs’ official names and their alias(es), country of origin, motives, first-seen year, activity patterns, and sponsors.
Collection of News Articles. Oftentimes, APTs occur in the context of cyber warfare. To identify plausible connections between APTs and external factors (including national conflicts, geopolitical events, and global crises) we gathered security news articles and media reports about APT incidents for the last decade. Note that our collection includes $177$ news articles (subsection 3.1) through the Google News search engine (News, 2024) using the APT and attack keywords.
Source Credibility. Characterizing the sources is crucial for further analysis since unreliable source(s) could severely distort our collected data, compromising the validity of the findings. First, we explicitly extract the origin of a report (e.g., organization name) from the collection of TRs. We confirmed that $1,412$ ($93.6\%$) of our TR collection come from highly reliable sources, including reputable security companies like Kaspersky (Lab, 2025a), RSA (Security, 2025c), FireEye (LLC, 2025), and Microsoft (Microsoft, 2025b), as well as governmental agencies like NATO (Organization, 2025), US-CERT ((2025), CISA), FBI (of Investigation, FBI), or trustworthy news sources. Figure 2 highlights the top $15$ sources, collectively representing approximately half of TRs. The remaining $97$ TRs ($6.4\%$) come from individual web blogs and security experts, which may be relatively less reputable or credible. Second, for the collection of TAs, subsection 3.1 shows that the original data comes from highly trusted companies and projects that deal with CTI, such as MITRE ATT&CK (Corporation, 2024b), MISP (Cynthia Wagner and Iklody, 2016), ETDA ((2019), ETDA), and CrowdStrike (CrowdStrike, 2025).
3.2. APT Information Retrieval
We carefully define ten items that have the potential to reveal longitudinal APT changes and global trends, including MITRE IDs (Corporation, 2024b), CVEs (Corporation, 2024a), YARA rules (VirusTotal, 2022), threat actors (APT groups), victim country, zero-day vulnerabilities, initial attack vectors, associated malware, target sectors, and attack durations.
3.2.1. LLM-based Retrieval
In this work, we leverage the language model’s contextual inference capabilities to accurately extract relevant information from our collected technical reports.
Table 4. Performance comparison of precision (P), recall (R), and F1 scores across three different LLMs. To evaluate LLM models, we obtain the ground truth by manually inspecting around 120 technical reports. Due to its superior performance, we choose the GPT-4-Turbo model (Section 3.2.1).
Language Model Gemini Flash (Anil et al., 2024) GPT-4-Turbo (OpenAI, 2024a) GPT-4o (Hurst et al., 2024)
Search Item P R F1 P R F1 P R F1
Threat Actor 0.98 0.78 0.87 0.98 0.80 0.88 0.98 0.77 0.86
Victim Country 0.84 0.77 0.80 0.88 0.86 0.86 0.82 0.77 0.79
Zero-day 0.96 0.65 0.77 0.95 0.95 0.95 0.89 0.74 0.81
Average 0.93 0.73 0.81 0.94 0.87 0.90 0.90 0.76 0.82
Table 5. Comparison of precision (P), recall (R), and F1 scores between a signature-based (e.g., IoCParser) and an LLM-based approach (e.g., GPT-4-Turbo). Note that IoCParser is capable of seeking CVE, MITRE ID, and YARA rules alone. (*) represents the items that we adopt GPT Turbo’s results that demonstrate the best LLM performance (Section 3.2.2).
Tool IoCParser (Tines, 2025) GPT-4-Turbo (OpenAI, 2024a)
Search Item P R F1 P R F1
CVE 0.98 0.95 0.97 0.97 0.84 0.90
MITRE ID 0.97 0.96 0.97 0.99 0.93 0.96
YARA 1.00 0.96 0.98 0.94 0.86 0.90
Attack vector* 0.89 0.77 0.83
Malware* 0.74 0.70 0.72
Target sector* 0.82 0.89 0.85
Average 0.98 0.96 0.97 0.89 0.83 0.86
LLM Prompt and Questions Design. We design our prompt and questions to enhance the accuracy of responses during the retrieval process of an LLM. We systematically evaluate varying question formulations and methodological techniques to identify the most effective strategies. Following the recommendations of Kumarasinghe et al. (Kumarasingh et al., 2024), we incorporate both the “Role Play” and “Specificity and Precision” means for prompt generation. The former approach enables the LLM to adopt a defined perspective, thereby generating more contextually appropriate and relevant responses, while the latter approach reduces the likelihood of irrelevant or inaccurate outputs. Likewise, for query construction, we employ “Specificity and Precision” to ensure simplicity and straightforwardness, which minimizes ambiguity and improves overall quality of the retrieved responses. Figure 9 and subsection A.1 in the Appendix are the final prompt and questions for our study.
LLM Model Choice and Evaluation. To obtain the desirable answers as reliably as possible, we sample three items for retrieval (e.g., threat actors, victim countries, zero-day vulnerabilities if any). To this end, we evaluate three popular LLMs: Gemini-1.5-Flash (Anil et al., 2024), GPT-4o (Hurst et al., 2024), and GPT-4-Turbo (OpenAI, 2024a) with the same prompt and questions (Figure 9 and subsection A.1 in Appendix). However, it is well-known that LLM models may generate inaccurate responses (i.e., hallucinations). This means we cannot completely trust the responses from an LLM. In response, we randomly picked around $120$ articles for manual inspection: i.e., a human compares the LLM responses with the ground truth in a technical report. Note that we use precision ($P=\frac{TP}{TP+FP}$), recall ($R=\frac{TP}{TP+FN}$), and F1 ($F1=\frac{2PR}{P+R}$) for evaluation metrics where TP, FP, and FN denote the number of true positives, false positives, and false negatives, respectively. Notably, FPs refer to cases where the LLM incorrectly identifies attributes that are not present in the report, while FNs arise when the model fails to detect attributes that are present. For interested readers, we present a selection of human-verified answers at subsection A.1 in Appendix. subsubsection 3.2.1 presents a performance comparison of three LLM models. GPT-4-Turbo achieves the highest scores in all Precision ($0.94$), Recall ($0.87$), and F1 score ($0.90$) over other LLM models.
3.2.2. Rule-based Retrieval
The IoCParser (Tines, 2025) tool is designed for processing IoCs from various data sources (e.g., CTI reports, security logs, other security-related texts). As the parser extracts three specific types of information, including CVEs, MITRE ATT&CK Technique IDs, and YARA rules, we compare it with GPT-4-Turbo. subsubsection 3.2.1 demonstrates that IoCParser surpasses GPT-4-Turbo in extracting CVEs, MITRE IDs, and YARA rules, achieving F1 scores of $0.97$, $0.97$, and $0.98$, respectively. Hence, we decided to include IoCParser’s results as a complementary tool. Our further examination of IoCParser’s retrieval failure reveals that 1 some TRs do not follow the standardized CVE format (e.g., year-vul_id instead of CVE-year-vul_id), which reduces the recall; 2 some TRs extracted from web pages have the parser retrieve irrelevant IoCs due to a noise in the extracted content, decreasing its precision.
Table 6. We investigate 10 items from each technical report on a specific APT. We adopt rule-based (e.g., IoCParser (Tines, 2025)) and LLM-based approaches. The number (ratio) of TRs denotes retrieved items out of 1,509 TRs because not every piece of information was available (Section 3.2.4).
Search Item Retrieval Approach # of TRs Ratio
CVE Rule 416 27.6%
MITRE ID Rule 175 11.6%
YARA Rule 131 8.7%
Threat actor LLM 1,089 72.2%
Victim country LLM 886 58.7%
Zero-day LLM 839 55.6%
Attack vector LLM 1,186 78.6%
Malware LLM 1,287 85.3%
Target sector LLM 1,228 81.4%
Attack duration LLM, Manual inspection 235 15.6%
As a final note, IoCParser does not have features to retrieve other items.
Figure 3. Summary of global APT trends over the past decade. Each box represents six key points for a given year: the number of APT campaigns, the number of associated malware samples, the most frequently attacked country, the primary threat actor, the most commonly used initial attack vector, and the most targeted sector. For the last 10 years, the most frequently targeted country, the most active APT group, the most predominant attack method, and most commonly targeted sector have been the United States, Lazarus group, malicious documents (and spear phishing almost equally contributed), and the government/defense sector, respectively (Section 4.1). Note that the numbers in parentheses represent the occurrences per year.
3.2.3. Manual Retrieval
We manually verify the accuracy of attack durations retrieved by the LLM to estimate both the lifecycle of individual APT campaigns and the time required to patch associated vulnerabilities. However, determining the precise lifecycle of an APT incident is inherently challenging due to its persistence and stealthy nature. Such campaigns may persist for weeks, months, or even years, depending on the adversary’s ultimate objective (e.g., disrupting target operations vs. stealing sensitive data). Hence, we define the start and end of an APT case based on the earliest and latest known discoveries of related activity, using day-level granularity. When exact dates are unavailable, we approximate using the midpoint of the month (i.e., the 15th day). Note that we exclude cases where the year-based (rough) information is solely available because such imprecision could lead to inaccurate estimations.
3.2.4. Availability of Retrieved Information
As one might imagine, not every piece of information was available in our collection of TRs. Indeed, we discover that only one TR contained all 10 items of interest. subsubsection 3.2.2 presents the number of TRs (out of $1,509$) that include each retrievable item. YARA rules and MITRE ID are sparsely available, appearing solely $8.7\%$ and $11.6\%$ of the TRs, respectively. In contrast, malware information and target sectors are prevalent, in approximately 8 out of 10 TRs ($85.3\%$ and $81.4\%$, respectively). Given that our analysis relies on the records in the collected TRs as a best-effort approach, we find it encouraging that a non-negligible number of items was successfully retrieved. We hypothesize that the covert nature of APT campaigns contributes to the limited availability of malware samples and lack of precise information regarding the duration of attacks.
3.2.5. Information Retrieval from Technical Reports
We convert web-based technical reports into PDF format using pdfkit (Golovanov, 2021) in conjunction with wkhtmltopdf (Truelsen and Kulkarni, 2022). Then, we utilize LangChain (LangChain, 2022) that offers a structured framework for data processing and retrieval. To enable context-aware question answering, we employ a Retrieval-Augmented Generation (Piktus et al., 2020) (RAG) pipeline by vectorizing the technical reports into a vector database. For each query, relevant passages are retrieved and incorporated into the prompt to provide contextually rich information to the LLM. In summary, our approach consists of the following phases: 1 extracting texts from each PDF document using PyPDFLoader (LangChain, 2025), 2 vectorizing those texts with the OpenAI’s (OpenAI, 2025) embedding model, 3 storing vectorized embeddings in the FAISS library (Johnson et al., 2019), and 4 generating responses by constructing prompt templates and chaining them with user queries to facilitate effective question answering. For performance comparisons (subsubsection 3.2.1), we access LLMs via LangChain’s API integrations, including GPT-4-Turbo (OpenAI, 2024a), GPT-4o (Hurst et al., 2024), and gemini-1.5-flash-latest (Anil et al., 2024). It is worth noting that achieving full accuracy and reliability remains challenging due to complications from PDF processing (Section 5).
3.3. Refining LLM-generated Responses
This section describes the refinement process of LLM-generated responses using our prompt and questions.
Normalization and Categorization. For our analysis, it is essential to interpret and normalize the responses generated by the LLM, particularly for the victim countries, attack vectors, and target sectors. To ensure consistency, victim countries were standardized by converting them into their corresponding two-letter country codes (for Standardization, 2025). For the categorization of attack vectors, we referred to the work of Sharma et al. (Amit Sharma and Saraswat, 2023), which identifies the most common attack vectors. This process results in classifying the attack vectors into the following 12 distinct categories: Spear Phishing, Phishing, Watering Hole, Credential Reuse, Social Engineering, Vulnerability Exploitation, Malicious Documents, Covert Channels, Drive-by Download, Removable Media, Website Equipping and Meta Data Monitoring. Meanwhile, for the categorization of target sectors, we adopt the taxonomy proposed in (Sectrio, 2024), which includes nine categories: Government and Defense Agencies, Corporations and Businesses, Education and Research Institutions, Critical Infrastructure (e.g., transportation, water supply), Financial Institutions, Individuals, Media and Entertainment Companies, Non-Governmental Organizations (NGOs) and Nonprofits, and Manufacturing. In line with (Google, 2025), we treat Energy and Utilities and Healthcare as distinct categories from critical infrastructure. Additionally, we include Cloud/IoT Services as a category to capture its relevance in recent attacks, such as those involving supply chains. This process results in classifying the target sectors into 12 distinct classes.
De-duplication and Filtering. After consolidating all sources from the collection of TRs and TAs, we identified a substantial number of duplicate entries. To address this, we extracted the full text from each PDF and computed cosine similarity using OpenAI’s embedding (OpenAI, 2024b). We apply an empirically-chosen similarity threshold of $0.85$. This reduces the number of TRs to $1,560$ unique dossiers after removing $1,003$ duplicates. Furthermore, we found that $51$ TRs focus on APT trends involving multiple campaigns rather than analyzing individual APT instances. We filter out these reports, resulting in a final dataset of $1,509$ TRs. Similarly, the number of TAs was reduced to $884$ unique APT groups after trimming $800$ duplicates. Next, we further filtered out the threat actors (n = $281$; 31.8%) that hold little information beyond their names, such as aliases and country of origin. Finally, we use $603$ identifiable APT groups for further analysis.
APT Name Aliases. The same APT group may operate under multiple aliases. For instance, the notorious APT group known as APT28 (Wikipedia, 2025f) is frequently referred to as FANCY BEAR, Pawn Storm, or Sofacy depending on the security vendor. To minimize complications, we de-duplicate APT groups with the following two-phase process. First, we use the identifiers as primary names from each threat actor source, along with their known aliases. Then, we merge alias entries across different sources based on these primary names. It is possible that despite these steps, due to the inconsistent naming conventions of APT groups across vendors and the absence of ground truth, some inaccuracies may remain in our final set.
3.4. Visual Representations
We develop the APT map (Figure 10 in Appendix) and its corresponding timeline chart using the React framework (Meta, 2025) and the amCharts library (amCharts, 2025) to enable interactive user experiences. The front-end has been integrated with a Flask-based backend (Pallets, 2024), deployed on the Heroku platform (Salesforce, 2025a). Additionally, we use Tableau (Salesforce, 2025b), a visual analytics platform, to construct a flow diagram illustrating the relationships between years, threat actors, and victim countries by incorporating the Sankey Viz extension (Tableau, 2024) (Figure 11 in Appendix).
4. Decadal Landscape of APT Campaigns
To longitudinally understand the landscape of APTs, we define the following research questions (RQs) under four themes.
  • RQ1: Evolution of APTs over a decade. How have APT campaigns evolved over the past 10 years regarding victim countries, threat actors, target sectors, initial attack vectors, and zero-day vulnerabilities?
  • RQ2: Cyber Threat Intelligence records for APTs. How is APT-related information captured across common threat intelligence sources, including vulnerability databases, attack frameworks, and indicators of compromise?
  • RQ3: Common traits of APTs. To what extent do APT campaigns exhibit concealment and aggressiveness?
  • RQ4: External dynamics affecting APTs. How do external factors, such as political events, international conflicts, global pandemics, or economic instabilities, affect APT activity?
(a) Number of APT attacks against the 10 most victimized countries over the past decade. The United States has experienced the highest number of attacks, followed by India, South Korea, Japan, and Russia.
(b) Number of APT campaigns conducted by the 10 most active threat actors over the past decade. The Lazarus group has been identified as the most active APT group, while APT28 is notable for its frequent exploitation of zero-day vulnerabilities.
Figure 4. Decadal trends in APT activity by victim countries (left) and threat actors (right). A circle size reflects the frequency of APT incidents, while color gradation represents the number of zero-day vulnerabilities associated with each entity as a concrete value (i.e., lighter red indicates fewer occurrences) (Section 4.1).
(a) APT campaigns consistently target the government and business sectors, as well as increasing attacks on the education and financial sectors in recent years.
(b) Malicious documents and spear phishing remain the most common initial attack vectors consistently harnessed by APTs, followed by vulnerability exploitation.
Figure 5. Decadal APT trends in 12 target sectors (left) and 12 initial attack vectors (right). We follow the categories of sectors from the guides (Sectrio, 2024; Google, 2025) and the attack vectors that Sharma et al. (Amit Sharma and Saraswat, 2023) proposed. The figures illustrate the distributions of each sector/vector over the last 10 years. The percentages within a stacked bar chart indicate the three most common target sectors and attack vectors for each year, along with their respective proportions. Note that a single APT case may entail multiple target sectors or attack vectors, which we count individually (Section 4.1).
4.1. Evolution of APT Campaigns Over a Decade
This section explores the temporal changes in APT campaigns. We use the end year of the attack identified by the LLM, since the start dates of APT operations are often unknown or partially known (around 27%) due to the covert nature of APT operations. Otherwise, we base our statistics on the year in which the corresponding technical reports were published. Accordingly, each APT is counted only once, even if it spans multiple years. However, we note that a single APT case may involve multiple target sectors or attack vectors, each of which is counted separately, as illustrated in Figure 5.
Comprehensive Overview. Figure 3 presents a comprehensive overview of APT trends over the past decade, highlighting key attributes for each year, including the most frequently targeted countries and sectors, the most active threat actors, and the most widely employed initial attack vectors. The United States (US) consistently appears as the primary target, while the Lazarus group (Wikipedia, 2025g) emerges as the most active threat actor during this period. Malicious documents constitute the most popular initial attack vector. Furthermore, the frequency of APT attacks involving malware closely aligns with the overall trend in APT activity, indicating that malware remains a core component of most campaigns.
Victim Countries. A total number of 154 countries were identified as victims across $1,509$ APT campaigns, representing around 80% of all nations worldwide. We analyze the 10 most victimized countries that account for $43.1\%$ ($650$) of all incidents. 3(a) presents the trends in APT attacks against these countries from 2014 to 2023. US remains the primary target throughout the decade with the exceptions of South Korea (in 2018, 2023) and Ukraine (in 2022). Despite minor fluctuations, our findings reveal that the following countries have also been heavily targeted: India (IN), South Korea (KR), Japan (JP), Russia (RU), China (CN), Great Britain (GB), Ukraine (UA), Germany (DE), and Türkiye (TR). A notable spike in attacks across the 10 most countries occurred in 2016 with $174$ recorded cases. There has been a steady decline in APT activity since 2021, with cases dropping from $157$ in 2020 to $121$ in 2021, and to $67$ in 2023.
Threat Actors. A total of $446$ unique threat actors have been identified in our APT dataset, $263$ of which are known to be state-sponsored (while others unknown). The 10 most active APT groups are responsible for 326 attack incidents, representing $21.6\%$ of all campaigns. 3(b) shows the number of attacks and the frequency of associated zero-day vulnerabilities by TAs. Among them, Lazarus (Wikipedia, 2025g), APT28 (Wikipedia, 2025f), and APT29 (Wikipedia, 2025c) stand out as the most prolific threat actors over the past decade, with $76$, $64$, and $35$ campaigns, respectively. APT28, in particular, exhibits a strong inclination toward exploiting zero-day vulnerabilities, with 22 instances accounting for $34.4\%$ of its operations. Interestingly, the activity levels of these threat actors are not uniformly distributed over time; rather, their campaigns tend to cluster within specific periods, likely reflecting shifting strategic objectives. For instance, the Sandworm (Wikipedia, 2025i) group conducted the majority of its operations within a concentrated five-year span. The peak year for activity among these top actors was 2017, during which $48$ attacks were recorded. Since 2021, however, a gradual decline in activity has been observed, with the number of attacks falling from $47$ in 2020 to $27$ in 2021, and further to $26$ in 2023.
Zero-Day Vulnerabilities. One of the common beliefs is that APT campaigns frequently exploit zero-day vulnerabilities due to their capability to bypass defenses and maintain stealth. We identify $204$ APT incidents ($13.5\%$) that incorporate zero-day vulnerabilities into their campaigns. 3(a) and 3(b) depict the annual use of zero-day vulnerabilities over the past decade from the perspectives of the top 10 victim countries and threat actors. Our analysis reveals notable peaks in zero-day exploitation between 2014 and 2016, with an average of $39.7$ cases affecting victim countries and $7.7$ cases linked to threat actors. The highest recorded number of zero-day incidents targeting countries occurred in 2015, reaching $50$ cases. In contrast, 2019 saw a sharp decline, with only nine cases affecting victim countries and just one attributed to threat actors. Since 2018, a general downward trend in zero-day vulnerabilities has been observed, possibly due to the increasing prevalence of 1-day vulnerabilities (i.e., non-patched systems) or rising costs and the complexity of developing zero-day vulnerabilities.
Target Sectors. 4(a) depicts the distribution of 12 sectors targeted by APT campaigns for 10 years, demonstrating notable shifts in attack patterns over time. The Government and Defense sector consistently accounts for a substantial share of APT activities, with an average of $65.6$ incidents per year. The second most targeted is the Corporations and Businesses sector, which has emerged as a prominent focus, averaging $48.5$ incidents annually. Notably, the third most frequently targeted sector varies across the years. The Education and Research sector ranks third in 6 out of 10 years, with a peak of $35$ cases in 2021 (possibly due to the COVID-19 pandemic). Meanwhile, the Critical Infrastructure sector has seen fluctuations, peaking $29$ in 2016 and declining since 2019.
Initial Attack Vectors. 4(b) illustrates the distribution of 12 initial attack vectors in APT attacks over the past decade. According to our investigation, malicious documents appear the most preferred initial access in APT campaigns, averaging $54.6$ incidents per year. The second most common attack vector is spear phishing, with an average of $53.6$ cases annually. This sophisticated phishing often entails sending fraudulent emails or deceptive messages to carefully selected individuals or organizations, rendering the attack more convincing. Malicious documents are often combined with spear phishing to gain the initial access to target system. Furthermore, across the whole decade, vulnerability exploitation consistently ranks as the third most frequently employed attack vector, with an average of $38.9$ incidents per year. This technique involves leveraging software vulnerabilities to gain unauthorized access to the system. Interestingly, watering hole attacks  [^4] A watering hole attack targets a specific group by compromising websites they frequently visit with malicious code. Upon successful exploitation, attackers can gain access target systems and exfiltrate information while maintaining long-term control., ranked as one of most utilized attack vectors in 2014, with $32$ instances, exhibited a gradual decline afterwards, with only $4$ cases recorded in 2023.
Key Takeaway 1: Over the past decade, APT campaigns have impacted 154 countries worldwide where US, IN, and KR remain frequently targeted. Although 446 unique threat actors have been identified, a small subset (Lazarus, APT28, APT29) accounts for a significant portion of attacks. Zero-day usage peaked between 2014 and 2016 but has declined in recent years. Government and corporate sectors stay the most targeted. Malicious documents and spear phishing dominate as initial penetration vectors.
4.2. CTI Records in APT Campaigns
This section explores how APT campaigns can be described across common threat intelligence sources, including a common vulnerability database, attack frameworks, and indicator of compromise.
Table 7. Top 10 most frequently observed MITRE ATT&CK techniques (Corporation, 2024b) in APT campaigns over the past decade. Notably, Execution, Defense Evasion, and Discovery are the most common tactic categories, using Command and Scripting Interpreter and Application Layer Protocol being among the most prevalent. These tactics are well aligned with the nature of APTs (Section 4.2).
MITRE ID Description Tactic Count Ratio
T1059 Command/scripting interpreter Execution 77 3.0%
T1071 Application layer protocol Command and control 76 2.9%
T1082 System information discovery Discovery 65 2.5%
T1027 Obfuscated files or information Defense evasion 60 2.3%
T1140 Deobfuscate/decode files or information Defense evasion 56 2.2%
T1041 Exfiltration over C2 channel Exfiltration 54 2.1%
T1204 User execution Execution 51 2.0%
T1053 Scheduled task/job Execution, persistence, privilege escalation 49 1.9%
T1083 File/directory discovery Discovery 47 1.8%
T1036 Masquerading Defense evasion 45 1.7%
MITRE IDs. We extract a total of $2,582$ MITRE ATT&CK techniques (Corporation, 2024b) from $175$ TRs available in our collection (11.6% in subsubsection 3.2.2), among which $263$ technique IDs are unique. Considering the whole $359$ distinct identifiers in the MITRE ATT&CK framework, this reveals that APT campaigns harness a wide spectrum of techniques to facilitate intrusions. subsection 4.2 demonstrates the 10 most frequently observed techniques ($22.5\%$ of all instances), along with their descriptions, tactics, occurrences, and proportional representation. As expected, the most commonly observed tactics in APT campaigns include 1 execution that involves the running code on local or remote systems; 2 defense evasion that aims to bypass detection mechanisms; and 3 discovery that focuses on reconnoitering intelligence about internal systems or networks. Additional frequently observed tactics are consistent with the advanced and stealthy characteristics of APT campaigns, such as command and control, exfiltration, persistence, and privilege escalation.
CVEs. We extract a total of $1,088$ CVEs from $416$ TRs available in our collection (27.6% in subsubsection 3.2.2), among which $431$ CVEs are unique. It is noted that multiple CVEs may be leveraged in a single APT campaign to achieve full-chain exploitation. Notably, the vulnerabilities exhibit high severity levels, with an average score of $8.5$. subsection 4.2 presents the top 10 most frequently observed CVEs ($22.8\%$ of all instances), along with their severity scores, vulnerability types, the number of affected software, occurrences, and proportions. Our further analysis shows that Microsoft Windows and Office are the most commonly targeted platforms/software, relating to $90\%$ of the identified CVEs. Unsurprisingly, the most prevalent vulnerability types include remote code execution (RCE), memory corruption, use-after-free (UAF), stack overflow and privilege escalation (PE). Notably, the CVE with the highest severity score ($9.3$), CVE-2015-5119, is associated with the exploitation of Adobe Flash Player, which was officially discontinued back in 2020.
YARA Rules. We extract a total of $419$ YARA rules from $131$ TRs available in our collection (8.7% in subsubsection 3.2.2), among which $419$ rules are unique. We hypothesize that the limited availability of YARA rules may be attributed to the sensitive nature of APT campaigns. Contributing factors include: 1 ongoing private investigations, 2 disclosure restrictions imposed by non-disclosure agreements or internal policies, 3 the risk of rule evasion or misuse by threat actors, and 4 high variability and obfuscation in malware samples, which complicates the creation of generalizable detection rules.
Key Takeaway 2: APT campaigns leverage a diverse set of intrusion techniques, with 2,582 MITRE ATT&CK instances, highlighting frequent use of tactics like execution, defense evasion, and discovery. Analysis of 1,088 CVEs reveals that Windows operating systems are the most targeted platforms, with remote code execution being the most common vulnerability type. The CVEs are highly severe, averaging the score of 8.5. On the other hand, YARA rules are scarce in public APT reports, likely due to confidentiality concerns, evasion risks, and the technical challenges posed by malware variability.
Table 8. Top 10 most frequently exploited vulnerabilities in APT campaigns. The APT groups show a strong preference for remote code execution (RCE) and memory corruption vulnerabilities, with CVE-2012-0158 being the most widely exploited (5.4%). High severity vulnerabilities such as privilege escalation (PE) and use-after-free (UAF) are also commonly leveraged. (Section 4.2). Note that we follow vulnerability naming from the National Vulnerability Database (Booth et al., 2013).
CVE Severity Vuln Affected S/W Count Ratio
CVE-2012-0158 8.8 (High) RCE 19 59 5.4%
CVE-2017-11882 7.8 (High) Memory Corruption 4 44 4.0%
CVE-2017-0199 7.8 (High) RCE 8 33 3.0%
CVE-2018-0802 7.8 (High) Memory Corruption 4 20 1.8%
CVE-2015-5119 9.8 (Critical) UAF 7 18 1.7%
CVE-2015-1641 7.8 (High) Memory Corruption 11 16 1.5%
CVE-2010-$33$ 7.8 (High) Stack Overflow 8 15 1.4%
CVE-2014-6332 9.3 (High) RCE 11 15 1.4%
CVE-2015-1701 7.8 (High) PE 3 15 1.4%
CVE-2014-4114 7.8 (High) RCE 10 13 1.2%
4.3. Common Traits of APT Campaigns
This section explores the underlying characteristics of APTs, mostly focusing on concealment and aggressiveness.
APT Duration.
Figure 6. CDF and histogram of APT campaign durations from 235 cases (subsubsection 3.2.2). The plot shows that around half of APT incidents lasted 137 days or fewer, while the remaining half extended beyond this duration. Notably, the longest recorded campaign persisted for 1,766 days, whereas the shortest lasted only a single day (Section 4.3).
Figure 6 presents the distribution of attack durations and their cumulative distribution function (CDF). Our findings indicate that approximately half of APT campaigns have lasted five months or less ($137$ days), while the remaining half have extended beyond that duration. We observe a significant variation in attack durations, ranging from a single day to nearly five years. The longest recorded campaign spans from June 2011 to April 2016 ($1,766$ days), which is associated with Project Sauron (GREAT, 2016). Attributed to a suspected US-based threat actor, this campaign primarily targets governmental and research institutions in Russia, Iran, Rwanda, and Italy. The second longest campaign lasts $1,706$ days, and was linked to the Iranian APT group Ajax Security Team as part of Operation Saffron Rose (Mandiant, 2014). This operation was first detected on July 12, 2009, and remained active until March 2014. At the other end of the spectrum, the shortest attack lasts only a single day, which targeted France’s TV5Monde broadcasting network (AhnLab, 2015). This incident, attributed to Cyber Caliphate that is reportedly linked to the Islamic State of Iraq and Syria, results in a shutdown on April 8, 2015. The second shortest attack, spanned just two days, involves the distribution of malware via Hangul document files (AhnLab, 2014). The campaign was observed between December 9 and 11, 2014, primarily targeting South Korea’s power infrastructure.
Vulnerabilities and Patches. We further investigate $62$ TRs that include both CVE identifiers and corresponding attack durations, recognizing $128$ distinct CVEs. We assume that the CVE release date corresponds to the availability of a patch. Figure 7 displays the relationship between the attack timelines and patch availability. Our findings reveal that around half of those CVEs (67 or $52.3\%$) are exploited as zero-day vulnerabilities, while the remaining half CVEs (61 or 47.7%) had been patched prior to the associated APT campaigns. On average, the time required to develop and release a patch for a zero-day vulnerability is approximately 200 days.
Figure 7. CVEs exploited in APT campaigns based on patch availability relative to the attack timeline. We consider a zero-day vulnerability when CVE (with a patch) becomes available during (shadowed region) or after APT campaigns. 61 one-day vulnerabilities (47.7%) were disclosed and patched prior to the APT campaign, whereas 67 zero-day vulnerabilities (52.3%) were exploited before a patch was available (Section 4.3).
Two-sided Nature as Both Attacker and Victim. APT campaigns reveal the dual role of nations, where a country can simultaneously act as both an aggressor and a victim. The APT campaigns in our dataset reveal that $23$ countries have been involved as attackers at least once, while $154$ countries have been targeted as victims in one or more incidents. We analyze the top 20 most attacking and victimized countries over the past decade, which account for $44.7\%$ of all APT cases. Notably, Figure 8 uncovers a significant asymmetry between attacker and victim nations. For instance, the CN-US attack ratio is 31:1, indicating that China has conducted $31$ times more attacks against the United States than the reverse. Similarly, both the KP-KR and RU-US attack ratios stand at around 15:1, reflecting substantial imbalances in cyber offensive activity. Russia, China, and Iran emerge as the most active attacking country, responsible for $229$, $232$, and $125$ APT cases (based on horizontal occurrences). Meanwhile, the United States, South Korea, and India stay the most frequently targeted, with $174$, $95$ and $89$ incidents. (based on vertical occurrences). Figure 8 also highlights the three most prominent bilateral attack relationships: namely, China targeting the US ($62$ cases), North Korea targeting South Korea ($46$ cases), and Russia targeting the US ($46$ cases). It is noteworthy to mention that not all the origin countries for APT groups are known.
Figure 8. Heatmap that depicts the distribution of APT campaigns based on the 20 origins (attacker country) and destinations (victim country), accounting for approximately 45% of the whole incidents over the past decade. The darker shades in each cell indicate the higher frequencies of APT cases. Notable patterns include a high volume of attacks originating from China, Russia, and North Korea, primarily targeting the United States, South Korea, and other regions. Notably, some countries are observed to engage in self-directed attacks (Section 4.3).
Self-directed APT Attacks. A closer examination of Figure 8 unveils that, in some APT campaigns, the origin and target countries are identical (i.e., main diagonal values of the matrix). For instance, Russia has been observed targeting its own systems in $29$ cases, while China appears to have done so in $21$ cases. Our further investigation identifies several contributing factors to these self-directed attacks. First, APT campaigns may target individuals within the same country, often focusing on political dissidents or human rights activists (Marczak et al., 2018; Ovi, 2023; Malhotra and McKay, 2021; International, 2021). These campaigns are typically conducted by domestic threat actors aligned with government interests. Second, such campaigns may also be directed at foreign companies operating within the country, particularly in sectors such as banking and payment systems (Carr, 2017; GREAT, 2019). Third, geopolitical and territorial disputes may drive this behavior, especially in politically tense regions such as the Russia–Ukraine border (Technologies, 2022; Milenkoski, 2023; Grunzweig, 2017; Falcone and Miller-Osborn, 2016). Finally, a few self-targeted incidents appear to result from unusual circumstances. For instance, the Longhorn threat actor –linked to the United States–compromised a US system, only to deploy an uninstaller within hours, suggesting the attack was likely unintentional (Johnson, 2017). Another notable case involves APT17 (FKIE, 2025), a China-associated group that reportedly targeted Chinese entities suspected of leaking sensitive information domestically (Intrusiontruth, 2019).
Key Takeaway 3: APT campaigns exhibit a wide range of durations, from a single day to nearly five years. About half of the 128 analyzed vulnerabilities are exploited as zero-days, while the rest are one-days. Notably, 23 countries have acted as attackers and victims in APT operations, revealing the dual-role nature of many nations and significant asymmetries (e.g., China launching 31 times more attacks on the United States than the opposite). Some campaigns were self-directed, with countries like Russia and China targeting domestic entities, often due to political repression or foreign company surveillance.
4.4. External Dynamics of APT Campaigns
Although APT operations are typically goal-driven and backed by sponsors, uncovering behind these attacks or external factors inherently remains a significant challenge. This section examines our curated collection of news articles and media reports from the web, discussing any known relationships (e.g., triggers, motives, or impacts) between global affairs and APT activities. We classify such external dynamics into four categories: political events, international conflicts, global pandemics, and economic gains.
Political Events. APT groups frequently exploit politically sensitive events to further their objectives. A notable example was the Russia-affiliated group Fancy Bear (APT28) (Wikipedia, 2025f), which launched a spear-phishing campaign targeting the Democratic National Committee (DNC) during the 2016 US presidential election (Kumar, 2016). As part of this operation, APT28 successfully breached systems linked to the presidential campaign, resulting in the theft of sensitive data. In a similar vein, Fancy Bear targeted Macron’s presidential campaign during the 2017 French election. The group infiltrated campaign systems, stealing credentials and sensitive data (Hern, 2017). Fancy Bear is also suspected of compromising the infrastructure of a German political party. By infiltrating the network, the group harvested critical information for political advantage (Guarnieri, 2015).
International Conflicts. APTs linked to international conflicts often involve the intersection of geolocational tensions and cyber warfare. For instance, during the Russo-Ukrainian War (Wikipedia, 2025h), the Russian-sponsored APT group known as Sandworm (Wikipedia, 2025i) hacked Ukrainian energy infrastructure (Wikipedia, 2025a) back in 2015, causing power outages that affected 230K consumers. Sandworm also attempted to disrupt the news agency in Ukraine (Vijayan, 2023) afterwards. Similarly, around 2023, the notorious Russian cyberespionage group, named Fancy Bear, launched an attack on a critical energy facility (Staff, 2023).
Global Pandemics. Beginning in late 2019, the COVID-19 pandemic brought the unprecedented impacts across various sectors worldwide (Wikipedia, 2025b). 4(a) shows a noticeable rise in APT attacks targeting the healthcare sector, with $14$ cases ($4.3\%$), and the education and research sector, with $33$ cases ($10.0\%$) in 2020, coinciding with the global spread of the virus. During this period, the China-sponsored APT41 (Wikipedia, 2025e) group exploited vulnerabilities in remote desktop services to attack healthcare organizations ((2020), HC3). Similarly, the Lazarus (Wikipedia, 2025g) group attempted to steal COVID-19-relevant intelligence (Parkk, 2020) by targeting a pharmaceutical company. Additionally, CozyBear (APT29) (Wikipedia, 2025c) has been suspected of attempting to steal the COVID-19 vaccines information (Bracken, 2020). The World Health Organization (WHO) was not spared, as the APT group called DarkHotel launched a password-stealing attack against WHO staff (Winder, 2020).
Economic Gains. APT activities are associated with the financial sector, often driven by economic gain. For instance, the Carbanak group (FIN7) (Wikipedia, 2024a) gained notoriety for its cyberattacks on Russian banking institutions in 2016 (Lab, 2016), coinciding with Russia’s gradual recovery from a prolonged economic recession (Group, 2016). With the rise of cryptocurrencies, many threat actors have shifted their focus to these digital assets as primary targets. Notably, the Lazarus group, reportedly sponsored by North Korea, was implicated to a major cryptocurrency theft in 2022 (BBC, 2022). Both FIN7 and Lazarus are recognized for their persistent targeting of financial institutions, underscoring the significant economic motivations behind their operations.
Key Takeaway 4: APT groups frequently exploit politically sensitive events, such as presidential elections, to infiltrate voting systems and influence public perception. Geopolitical conflicts often coincide with cyber operations, with Russian APTs launching attacks during international tensions. During the COVID-19 pandemic, healthcare and research sectors experienced a spike in APT campaigns (e.g., APT41, Lazarus, APT29), targeting pandemic-related organizations. Economic motives also drive APT activity (e.g., FIN7, Lazarus), conducting cyberattacks on financial institutions and cryptocurrency platforms reflecting a persistent focus on financial gain.
4.5. Visual Representations
We design an interactive map to visualize APT campaigns, enabling users to explore detailed information by selecting either an attacking or victimized country. The map integrates decade-long historical data by year, presenting key attributes such as associated threat actors, source(s), CVE identifier(s), initial attack vector(s), related malware, targeted sector(s), and estimated attack duration (when such information is available in the corresponding technical report). More importantly, the map maintains up-to-date information by dynamically retrieving content from technical reports using an LLM. Note that the source has been currently linked to TR#1 (CyberMonitor, 2024). In addition, we incorporate a timeline chart that links APT campaigns to relevant news articles for contextual reference. Figure 10 in Appendix presents a sample view of an APT campaign from 2020, in which APT41 (Wikipedia, 2025e) targeted Canada. Finally, we display an additional interactive diagram (Figure 11 in Appendix) that visualizes the relationships between the top 10 threat actors and the 30 most frequently targeted countries over the past decade.
5. Discussion and Limitations
This section discusses the limitations of our work.
Representativeness of APT Campaigns. Although our dataset covers a wide spectrum of APT campaigns, recent incidents likely remain unreported or undocumented. For instance, publicly available technical reports in recent years appear less comprehensive, as shown by the decline observed in 2023 (subsection 3.1). Besides, due to the covert nature of APT operations, capturing every case is inherently infeasible. While we chose three independent repositories to include as many APT instances as possible, community-aggregated datasets may still introduce selection biases and coverage gaps. Nonetheless, we believe that the collective intelligence drawn from both official and unofficial sources worldwide provides a representative sample, which would be sufficient to approximate a meaningful reflection of the broader ground truth.
Limited Responses from an LLM. While extracting targeted information using an LLM can be highly effective, it comes with several limitations related to accuracy, reliability, and completeness. First, PDF files are unstructured, leading to extract misordered or misaligned content during extraction. Complex layouts including tables and figures could further result in broken or fragmented text. Second, long documents must be segmented into smaller chunks due to the limited context window of LLMs, which may disrupt contextual coherence. Third, LLMs are susceptible to hallucination, potentially creating inaccurate or fabricated information when encountering incomplete or ambiguous input. Lastly, extracted text may contain extraneous or irrelevant content such as page numbers, headers, footers, legal disclaimers, or noises. With the possibly above reasons, our empirical assessments with sampled TRs (subsubsection 3.2.1) demonstrate that the GPT-4-Turbo model achieves around $90\%$ F1 scores. Nevertheless, we acknowledge that our study represents a best-effort approach, as the retrieval process is inherently dependent on the interpretive capabilities of the LLM.
Attack Duration. We understand that estimating the attack duration of APT campaigns is inherently challenging due to the covert and persistent nature of these operations. First, the reports may have inconsistent timelines across multiple reports by different organizations. Second, the reported dates may be distorted due to the unwillingness of a victim or national security reasons. Third, reliance on public sources may miss internal or classified timelines. Lastly, inferring the end date may be open because detecting the last known activity is difficult to determine. This work attempts to reconstruct durations based on the (known) records.
CVE and Patch Timing. In many cases, a patch is available at the time of CVE publication. However, the timing of a CVE’s release does not always guarantee patch availability, as it relies on responsible disclosure practices. Note that our study aims to approximate the timeline required to develop patches for zero-day vulnerabilities exploited in APT campaigns.
Future Work. While we carefully examine attack durations (Section 4.3) and attacker motivations (Section 4.4), there remain additional opportunities to gain deeper insights. One promising direction is to further investigate the relationship between attack duration and attacker objectives. Another is to uncover the evolution of persistent techniques and remediation trends over time.
6. Related Work
We categorize prior APT works into three main areas, including the detection and evaluation, Cyber Threat Intelligence, dataset regarding APT campaigns.
APT Detection and Evaluation. A substantial body of research has focused on the detection and evaluation of APT campaigns, including the development of frameworks such as HOLMES (Milajerdi et al., 2019), ProvG-Searcher (Enes Altinisik, 2023), Zimba et al. (Aaron Zimba and Chishimba, 2020), Marchetti et al. (Mirco Marchetti and Guido, 2016), and Hassan et al. (Wajih Ul Hassan and Marino, 2020). Notably, MAGIC (Jia et al., 2024) leverages graph neural networks and self-supervised learning to construct behavior graphs from system logs, applying a masked graph learning strategy to capture relationships between entities and events. On the other hand, CAPTAIN (Wang et al., 2025) introduces a rule-based intrusion detection system, which enhances traditional provenance-based approaches by learning fine-grained detection rules through gradient descent optimization. Wang et al. (Wang et al., 2024) propose a provenance graph-based detection framework by reconstructing attack chains and countering adversarial strategies through empirical evaluation. Meanwhile, Shen et al. (Shen et al., 2024) further contribute to the field by systematically evaluating the effectiveness of modern Endpoint Detection and Response (EDR) systems in detecting and mitigating adversarial tactics commonly used by APT groups in realistic attack scenarios. Malik et al. (Vaibhav Malik and Nalluri, 2024) present a multi-layered mitigation framework, evaluating network-based, host-based, and AI-driven detection methods. Our work differs from previous approaches, with a focus on trends and insights over the last decade from a macroscopic perspective (e.g., temporal and global analysis).
Cyber Threat Intelligence. Several studies have concentrated on the collection and evaluation of CTI. With an in-depth analysis of 22 APT reports, Ussath et al. (Martin Ussath and Meinel, 2016) identify common tactics, tools across different stages of APT campaigns. Similarly, Bahrami et al. (Bahrami et al., 2019) propose a taxonomy based on the cyber kill chain model to systematically categorize the tactics, techniques, and procedures (i.e., TTPs) by threat actors. Kumarasinghe et al. (Kumarasingh et al., 2024) present a multi-stage ranking framework designed to identify and prioritize the most relevant MITRE ATT&CK techniques by leveraging a combination of pre-trained and fine-tuned language models. Meanwhile, TTPHunter (Nanda Rani and Shukla, 2023) introduce an automated system that applies machine-learning-based sentence classification to extract MITRE ATT&CK-aligned TTPs from unstructured APT reports. Note that our work uses an LLM to retrieve pre-defined questions from TRs.
APT Dataset. TREC (Lv et al., 2024) collects a dataset by capturing APT behaviors at the kernel level, making it publicly available. Similarly, Kumarasinghe et al. (Kumarasingh et al., 2024) introduce an open benchmark dataset for training and validating varying techniques used in APT campaigns, which contains threat behavior descriptions from real-world reports and MITRE ATT&CK technique identifiers. Siracusano et al. (Siracusano et al., 2023) provide a collection of publicly available technical reports and corresponding CTI information. Moreover, Stojanović et al. (Branka Stojanović and Kleb, 2020) provide a comprehensive review of existing APT datasets and the frameworks employed to model attacks for the development of automated detection techniques. Their analysis highlights the limited availability of publicly accessible datasets and discusses the considerable challenges involved in collecting realistic, high-quality attack data. In this work, we compile several (large) collections of prior technical reports (e.g., 2,563 before dataset refinement), threat actor profiles, and related news articles.
7. Conclusion
This study presents a comprehensive, decade-long (from 2014 to 2023) analysis of APT campaigns, offering a macroscopic perspective of how these threats have evolved across countries, sectors, and attack techniques. By leveraging a hybrid information retrieval approach that combines LLM inference and rule-based extraction, we systematically process and analyze over 1,500 technical reports, revealing key trends in threat actor behavior, vulnerability exploitation, and campaign duration. Our key findings highlight the global reach, persistent nature, and strategic targeting patterns of APT groups, while also uncovering the contextual influence of geopolitical events, global crises, and economic motivations. Although most observations align with our expectations, our findings also discover interesting (but less known) facts such as a wide range of attack durations, self-directed attacks, notable prevalence of one-day CVEs. To support ongoing research and awareness in the area of APTs, we release an interactive visualization platform, which promotes deeper engagement with the evolving threat landscape.
Acknowledgments
We thank the anonymous reviewers for their constructive feedback. This work was partially supported by the grants from Institute of Information & communications Technology Planning & Evaluation (IITP), funded by the Korean government (MSIT; Ministry of Science and ICT): No. RS-2022-II221199, No. RS-2024-00437306, No. RS-2024-00337414, No. RS-2025-25457342, and No. RS-2025-25394739. Additional support was provided by the National Science Foundation under awards CNS-2126654, CNS-2440819, and DGE-2335798. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsor.
Appendix A Appendix
A.1. Notable APT Groups
Among the threat actors, we examine the 10 most active APT groups over the past 10 years. The Equation Group (Lab, 2015) is a highly sophisticated threat actor suspected of being tied to the USA. APT28 (Wikipedia, 2025f), affiliated with Russian intelligence, is known for its primary targets in the government, military, and security sectors. Lazarus Group (Sanger and Perlroth, 2014) is allegedly run by the government of North Korea. They have been responsible for several cyberattacks since 2010. APT41 (Wikipedia, 2025e), believed to operate under Chinese state-sponsored initiatives, is notorious for its dual-purpose operations, conducting both espionage and financially motivated campaigns. Turla (ATT&CK, 2024c), known as the Russian-linked cyber-espionage group, is recognized for its stealthy and sophisticated operations targeting government and diplomatic entities. APT34 (ATT&CK, 2024b), suspected to be associated with Iran, focuses on cyber-espionage against government, financial, and energy sectors. APT29 (National Cyber Security Centre, NCSC), also linked to Russia, targeted Western governments and scientific institutions, particularly during the COVID-19 pandemic, in attempts to obtain vaccine-related data. FIN7 (Wikipedia, 2024a), a financially motivated group with ties to Russia, has conducted large-scale cyber crimes against the hospitality and retail sectors, stealing massive amounts of payment card data. The MuddyWater group (ATT&CK, 2024a), suspected of being linked to Iran’s government, has carried out cyber-espionage campaigns targeting government and private organizations across the Middle East, Europe, and North America. Sandworm (Wikipedia, 2025i), connected to Russian intelligence, is infamous for its disruptive attacks on critical infrastructure, including power grids and industrial control systems. Finally, APT10 (Wikipedia, 2024b) is a Chinese state-sponsored cyber-attack group, mainly targeting aerospace, engineering, and telecom firms.
Table 9. Questions to obtain the responses from LLM.
Item Query
Threat Actor “What is the name of a threat actor group?”
Victim country “Which countries are being targeted?”
Zero-day “Was a zero-day vulnerability used in this attack? Answer with TRUE or FALSE.”
Attack vector “What are the initial attack vectors described in this report? Group them into one of followings: Spear Phishing, Phishing, Watering Hole, Credential Reuse, Social Engineering, Exploit Vulnerability, Malicious Documents, Covert Channels, Drive-by Download, Removable Media, Website Equipping, Meta Data Monitoring.”
Malware “Which specific malware, tool names, of software frameworks are used in the attack from this report?”
Attack duration “Analyze the following dates extracted from an article discussing an APT attack. Identify the start and end dates of the attack timeline, considering the dates provided may indicate various stages of the attack. Return the start date as the earliest reference to attack activity and the end date as the latest reference of activity specific to the attack. All Dates from article: [ALL_DATES_FROM_ARTICLE]. Additionally, the article’s publication date is provided for context: [PUBLISH_DATE].”
Target sector “Identify the targeted sectors in this document. Group them into one of followings: Government and defense agencies, Corporations and Businesses, Financial institutions, Healthcare, Energy and utilities, Cloud/IoT services, Manufacturing, Education and research institutions, Media and entertainment companies, Critical infrastructure, Non-Governmental Organizations (NGOs) and Nonprofits, Individuals.”
Label Initial Attack Vector Label Target Sector
V1 Spear-Phishing S1 Government and Defense
V2 Phishing S2 Corporations and Businesses
V3 Watering Hole S3 Critical
V4 Credential Reuse S4 Financial
V5 Social Engineering S5 Healthcare
V6 Vulnerability Exploitation S6 Energy and utilities
V7 Malicious Documents S7 Individuals
V8 Covert Channels S8 Education and Research
V9 Drive-by Download S9 Media and Entertainment
V10 Removable Media S10 NGOs and Nonprofits
V11 Website Equipping S11 Manufacturing
V12 Meta Data Monitoring S12 Cloud and IoT services
Table 10. Labels for initial attack vectors and target sectors.
You are an experienced security engineer analyzing the security articles describing cases of APT attacks. You need to answer the questions in the scope of the provided file as precise, accurate and short as possible. For each question just give me the answers straight without additional explanation of it. If the information is not mentioned answer with “Not mentioned”. Given below are the contents of the file and question of the user. {”context”} = [YOUR_CONTEXT] {”question”} = [YOUR_QUESTION]
Figure 9. LLM prompt to retrieve desirable information.
Figure 10. Snapshot of our APT map with victim countries around 2023. A total of 208 attacks and 81 involving zero-day vulnerabilities were identified that year. Selecting a country (e.g., Canada) in the map reveals associated threat actors, and choosing a specific actor (e.g., APT41 (Wikipedia, 2025e)) displays the details of the corresponding APT campaign below such as CVE information, malware, zero-day vulnerability, and initial attack vector (Section 4.5).
Figure 11. Flow diagram that depicts the relationships between the top 10 threat actors and the top 30 victimized countries over the past decade. Three columns represent the year, threat actor and victim country, respectively (Section 4.5).
Table 11. Selected APT campaigns along with corresponding information, including a threat actor(s), victim country(ies), zero day (ZD) vulnerabilities, associated malware, attack duration, initial attack vector(s), and target sector(s). The curated set represents the samples used for manual inspection to evaluate the accuracy of LLM’s repsonses with retrieval. The complete list of label definitions (e.g., V1-12, S1-12) is provided in subsection A.1. Note that a dash (-) indicates missing or unavailable information.
Year Threat Actor Victim Countries ZD Malware or Tools Attack Duration Attack Vectors Target Sectors
2014 Pawn Storm US, HU, AT, DE, FR, IQ, PL, PK SEDNIT V1, V2 S1, S2, S9
2014 Quedagh UA, GE BlackEnergy Apr ’13 – May ’14 V5, V7 S1
2014 Deep Panda US, JP, MY, HK Derusbi trojan, implants and droppers, RATs, DLL backdoor V1, V5 S1, S2, S3, S6, S10
2014 NEWSCASTER US, IL, UK, SA, SY, IQ, AF IRC Bots, data profiling tools V2, V5 S1, S2, S7
2015 Naikon APT PH, MY, KH, ID, VN, MM, SG, TH, NP, LA Remote administration utility, keyloggers, WinPcap network sniffers, RTLO, spyware V1 S1, S2
2015 Pawn Storm US, CA Keyloggers, C&C servers, sandbox V1 S1, S2, S7
2015 Butterfly BR, CN, HK, IN, IL, JP, KZ, MY, MA, NG, TW, TH, KR, AE Jiripbot, PintSized, Hacktools Apr ’12 - Jun ’15 V3, V6 S1, S2, S3, S6
2015 IL, PS DownExecute, Xtreme RAT, Poison Ivy Jun ’14 - Apr ’15 V1 S1
2016 Lazarus KR OnionDog, Icefog backdoor, USB worms, dropper trojans V1, V7, V10 S3, S6
2016 Scarlet Mimic RU, IN FakeM RAT, MobileOrder, CallMe, SFX archives V1, V7, V6, V3 S1, S7, S10
2016 NetTraveler RU, MN, BY, TR, UA NetTraveler, Saker, DarkStRat, Gh0st RAT, Netbot, PlugX V1, V6, V7 S1, S2, S7, S10
2016 StrongPity IT, BE, DZ, CI, MA, FR, TN, NL, CA, TR StrongPity, trojanized WinRAR, TrueCrypt Dec ’15 - Sep ’16 V3, V11 S3, S7
2017 Lazarus PL Shellcode, encrypted EXE files, cambio, perfmon Oct ’16 - Feb ’17 V8 S4
2017 APT28 US DealersChoice framework V1, V6, V7 S1, S2
2017 APT34 POWRUNER, BONDUPDATER V1, V6, V7 S1
2017 NewsBeef SA Shamoon, StoneDrill Nov ’16 - Jan ’17 V1, V4, V5 S2, S3
2018 OilRig OopsIE Trojan, QUADAGENT Jul ’18 - Sep ’18 V1, V7 S1
2018 Goblin Panda RU Sisfader RAT V7 S1, S2
2018 Patchwork IN, PK BADNEWS, EPS Dec ’17 - Mar ’18 V6 S1, S3
2018 Lazarus Fastcash, EternalBlue V6 S4
2019 OceanLotus VN, KH, TH, LA, CN OceanLotus RAT, Dll hijack, macOS backdoors, CocCocUpdate dropper Apr ’12 - Mar ’19 V2, V6, V7 S1, S2, S4, S8, S9
2019 APT28 KZ Zebrocy V2, V7 S2
2019 APT39 US, EG, ES, SA SEAWEED, CACHEMONEY, POWBAT, ANTAK, ASPXSpy, Mimikatz, BLUETORCH, REDTRIP, PINKTRIP, BLUETRIP Nov ’14 - Dec ’18 V1, V4, V6, V7 S1, S2, S3
2019 Lazarus RU KEYMARBLE, VBScript, CAB V7 S2, S3
2020 AridViper IL Augury malware addon Nov ’20 - Dec ’20 V2, V5, V7 S1
2020 Transparent Tribe IN Crimson RAT, Oblique RAT, AhMyth android RAT V2, V5, V7 S1, S7
2020 Darkhotel CN, JP Ramsay V6, V7, V11 S1, S3
2020 Kimsuky KR DLL hijack V1, V7 S1, S7
2021 Earth Wendigo TW JavaScript and WebSocket backdoors, XSS attack scripts, service worker script, WENDIGOE Trojan, Cobalt Strike V1, V8, V11 S1, S7, S8
2021 NAIKON RainyDay, Nebulae, Aria-body Jun ’19 - Mar ’21 V6, V8 S1
2021 GhostEmperor Jul ’20 - Jul ’21 V6 S1, S2
2021 Water Kappa JP Cinobi, SHELLOAD V9, V11 S4, S7
2022 Bitter APT BD ZxxZ, Almond RAT Oct ’21 - May ’22 V6, V7 S1
2022 Molerats APT PS, TR Spark, .NET-based backdoor, malicious RAR files Jul ’21 - Dec ’21 V1, V7 S1, S4, S9, S10
2022 APT-Q-43 PK VajraSpy RAT Jun ’21 - Feb ’22 V5, V9 S1, S7
2022 Lazarus KR PowerShell scripts Mar ’22 - Apr ’22 V1, V6, V7 S2, S3, S4, S7
2023 APT36 IN ElizaRAT, LimePad V9, V11 S1, S7
2023 AeroBlade US MS Office documents Sep ’22 - Jul ’23 V1, V7 S2
2023 APT28 DE, NL, UK, US, UA Jaguar Tooth V6 S1, S2
2023 APT-C-36 CO, EC Quasar RAT, Meterpreter V1, V7, V9 S1, S2, S7